OpenStudy (anonymous):

Solve this system of linear equations. 9x=1-4y -7x=-7+4y

OpenStudy (anonymous):

x = -3 y = 7

OpenStudy (anonymous):

add both the equations

OpenStudy (bahrom7893):

Here are the steps: 4y = 1-9x -7x = -7 + 1 - 9x 9x-7x = -6 2x = -6 x = -3 4y = 1-9*(-3) 4y = 1+27 4y = 28 y=7